OpenStudy (anonymous):

(3n^2-n) / (n^2-1) / (n+1)/n^2

OpenStudy (anonymous):

\[\frac{ 3n ^{2}-n }{ n ^{2}-1 }\div \frac{ n+1 }{ n ^{2} }\]

OpenStudy (anonymous):

factorising here, is the "key" to your solution.

OpenStudy (anonymous):

\[\frac{ n(3n-1) }{(n-1)(n+1) }\ \div \frac{ n ^{2} }{ n+1 }\] is this right

OpenStudy (anonymous):

YOu're on the right track. Now flip the second fraction upside down so that you can change the division sign to a multiplication sign.

OpenStudy (anonymous):

You already flipped it.

OpenStudy (anonymous):

sorry I was suppose to put multiply instead of divide

OpenStudy (anonymous):

Changed the sign. Okay.

OpenStudy (anonymous):

Then take away those common multiples.

OpenStudy (anonymous):

Wait I don't think you can take any.

OpenStudy (anonymous):

You sure you wrote down the correct equation. Because you can't take away any common values.

OpenStudy (anonymous):

Ya that is what they gave me

OpenStudy (anonymous):

Was your question: (3n^2-n) / (n^2-1) / (n+1)/n^2 meant to be this:(3n^2-n) / (n^2-1) / n^2/n+1?

OpenStudy (anonymous):

look at the last fraction, that does seem dodgy to me. Could you double check your fractions to make sure each term is in the correct position.

OpenStudy (anonymous):

\[\frac{ n ^{2} }{ n+1 }\] this is the original

OpenStudy (anonymous):

Ahah, I want to say told you so, but I won't. So now you flip that upside down. Now you can get rid of the common factors diagonally across each fraction.

OpenStudy (anonymous):

\[\frac{ n(3n-1) }{ (n-1)(n+1) }\times \frac{ n+1 }{ n^2 }=\frac{ 3n-1 }{ n-1 }\times \frac{ 1 }{ n }\]

OpenStudy (anonymous):

That's what you would end up with, correct?

OpenStudy (anonymous):

sorry I wrote it down and then flipped it back around.

OpenStudy (anonymous):

No worries.And then you just combine the result to leave it in one fraction.

OpenStudy (anonymous):

\[\frac{ 3n-1 }{ n-1 }\times \frac{ 1 }{ n }=?\]

OpenStudy (anonymous):

What you fill in the question mark, is your answer.

OpenStudy (anonymous):

\[\frac{ 3n-1 }{ n(n-1) }\]

OpenStudy (anonymous):

You nailed it right into the head. Well Done. Excellent work.

OpenStudy (anonymous):

Thank you for walking me through it.

OpenStudy (anonymous):

No worries, anytime.